An investment consultant has deposited $5000 of a clients money into a 10% simple interes account. How much should be deposited into an 18% annual simple interest account so that the effective annual interest rate on both deposits is 14%
Answer by stanbon(48568)

Let the amount invested at 18% be "x".
EQUATION:
interest + interext = interest
0.10(5000) + 0.18x =0.14(5000+x)
Multiply thru by 100 to get:
10(5000)+18x = 14(5000+x)
50000 + 18x = 70000+14x
4x=20000
x=$5000.00 (amount invested at 18%)
Cheers,
Stan H.